OpenStudy (asad786):

plz help What is the percent of decrease from 320 to 32 A. 10% B. 90% C. 100% D. 288%

OpenStudy (mathmale):

What have you done so far? What questions in your mind has this problem generated? We will compare the decrease in the original 320 to 320 itself: decrease in 320 to 32 ------------------- * 100% 320

OpenStudy (mathmale):

What is the decrease in 320 if we end up with 32? Explain how to find this decrease.

OpenStudy (decarr432):

Hi @mathmale would it be appropriate if i try and help?

OpenStudy (mathmale):

Oh, definitely, and thank you for the offer and for being a gentleman!

OpenStudy (mathstudent55):

There is a simple formula for percent change. If the percent change is positive, it is a percent increase. If the percent change is negative, it is a percent decrease. Look up the formula and use it.

OpenStudy (decarr432):

Okay @Asad786 The way I like to do percents isnt the best way but it works for me take your original nuumber say 320 and multiply by one of your answer choices for this example i wont ill use 50% and what i get for an increase if 160 and i know then that, it wouldn't work for the answer

OpenStudy (decarr432):

How about that @mathmale seem good?

OpenStudy (mathstudent55):

\(percent~change = \dfrac{new ~amount - original~amount}{original ~amount} \times 100\%\)

OpenStudy (mathstudent55):

In your case, the number went from 320 to 32, so the new amount is 32, and the old amount is 320. Plug in those numbers in the formula, and tell me what you got.

OpenStudy (tommynaut):

Rote-learning a formula works, but I think it's better to first understand the concept of percentages, similar to how @Decarr432 was suggesting. I first like to start at 100%; I know that this is the whole of the amount, so 100% of 320 is 320. If I was to get rid of 100% of 320, I'm getting rid of the whole thing (320 - 320)! If I'm adding 100% of 320, I'm effectively doubling it (320 + 320). In your question, you're given that it's a percent decrease. This means you'll be doing a subtraction, so 320 minus something. In fact, it would fit the equation 320 - x = 32. You can solve this equation to find x (no spoilers here) and then work out what x/320 is as a percentage, and that's your percentage change. Alternatively, maybe you've realised that 32 is 10% of 320; that gives a strong hint as to what the percentage decrease would be (how much of 100% do you get rid of to end up with 10%?). Another alternative would be to eliminate the answers you think can't possibly make sense (although you'd need to understand the concepts anyway if you were taking this route).

jhonyy9 (jhonyy9):

- so i think the best easy way us the rule of three simple 320 --------- 100% 32 --------- x ------------------ 32*100 x= ---------- = ? % 320 - so after this just subtract from 100% the result of x and will get the right answer on your question